OpenStudy (mendicant_bias):

How exactly do you take the complex conjugate of a term with an imaginary unit in the exponent? (Example below.)

OpenStudy (mendicant_bias):

If I was trying to take the complex conjugate of an expression like\[e^{\alpha + i \beta}\]How would I get to the result? I see countless examples of the complex conjugate taken on the web where there is simply a single term in the exponent, like \[e^{i \beta}\]where the complex conjugate is \[e^{-i \beta}\] Which makes sense enough, but how would I arrive at the result for the first exponent? I'm really hesitant to just mirror the rules applied to a normal complex number like taking the complex conjugate of a+bi, *what rules do I follow* by to get to the result when it's a more complex term than the second? @agent0smith

OpenStudy (mendicant_bias):

(maybe?) figured it out here: http://www.math.pitt.edu/~sparling/23012/23012complex1/node13.html Let's see.

OpenStudy (mendicant_bias):

Oh, whoops. Wrong section.

OpenStudy (agent0smith):

Well \[\large e^{a+ib} = e^a*e^{ib}\] so I'd think you just change the sign on the i term to get the conjugate.

OpenStudy (kainui):

There are a few ways you could do this depending on what sorta thing you want. Personally I can see a proof and be dissatisfied so I'd rather feel convinced in a way that makes it common sense to me. So, that being said, one semi lame way that might convince you is: \[(e^{a+bi})^* = \left(\sum_{n=0}^\infty \frac{(a+bi)^n}{n!} \right)^* =\sum_{n=0}^\infty \frac{(a-bi)^n}{n!} = e^{a-bi}\] But maybe something else is better idk how you feel about this one haha.

OpenStudy (anonymous):

Why not just use Euler's formula? (Maybe it's listed in the link above, I didn't check.) If you're okay with \(e^{it}=\cos t+i\sin t\), then \[\begin{align*}\overline{e^{x+iy}}&=\overline{e^x}\overline{e^{iy}}\\[1ex]&=e^x\overline{(\cos y+i\sin y)}\\[1ex]&=e^x\left(\overline{\cos y}+\overline{i\sin y}\right)\\[1ex]&=e^x(\cos y-i\sin y)\\[1ex]&=e^xe^{-iy}\\[1ex]&=e^{x-iy}\end{align*}\]